OpenStudy (314dokg):

pls guys. help me solve for x... just give me process by process. iforgot how to. and its confusing me. so pls 5(360/x)-9 = (x-2)180 / x

OpenStudy (amistre64):

multiply thru by x

OpenStudy (314dokg):

meaning to say i have to do lcd kind of thing? multiply both side by x?

OpenStudy (amistre64):

you have an x inthe denominator on both sides, to clear the fractions i would multiply thru by x; yes

OpenStudy (314dokg):

so ill get. 5x (360) - 9 = (x-2)180 is that how it works?

OpenStudy (amistre64):

almost, but i see that my reading of the setup was a little off 5(360) - 9/x = (x-2)180 is what you should get; that 9 was mean to me ... ive got a simple fix tho

OpenStudy (314dokg):

so meaning to say. that 5(360/x) are in one 1 fraction. while 9 isn't so i hav to multiply 9 by x instead of 5

OpenStudy (amistre64):

youre actually doing better at this than I am this morning :)

OpenStudy (amistre64):

5(360) - 9x = (x-2)180

OpenStudy (314dokg):

so it will look like this? \[5(360) - 9x = (x-2) 180\]

OpenStudy (amistre64):

yes, now you can distribute the 180 thru if need be

OpenStudy (314dokg):

thanks alot :D now i can do the other number easily :D

OpenStudy (amistre64):

you did great :)

OpenStudy (314dokg):

dude. theres a problem

OpenStudy (amistre64):

5(360/x)-9 = (x-2)180 / x 5(360)-9x = (x-2)180 5(360)-9x = 180x-2(180) 5(360)+2(180) = 180x+9x 12(180) = 189x 12(180)/189 = x is about 11.42 in decimal form; or just keep it in reduced fraction form

OpenStudy (314dokg):

dude. i post a new one with the problem :) just check it out. thanks :)

OpenStudy (amistre64):

12(180)/189 4(180)/63 4(60)/21 4(20)/7 = 80/7